1. A group consists of one hundred and forty students. Of these, eighty are studying Norwegian, forty are studying Swedish and fifteen are studying Norwegian and Swedish. A student is selected at random. What is the probability that a student is studying Norwegian or Swedish?
Answer #1

n(A\bigcap B)=n(A)+n(B)-n(A\bigcup B)

By using above formulae

n(A)=80

n(B)=40

n(A\bigcap B)=15

so we apply above formulae

15=80+40-n(A \bigcup B)

n(A \bigcup B)=80+40-15

=105

so answer is 105